OpenStudy (anonymous):

Solve the equation for y. 8x - 9y = 11 A. Y = -9/8x - 9/11 B. Y = 9/8x + 11/8 C. Y = -8x - 11 D. Y = 8/9x - 11/9

OpenStudy (anonymous):

8x-9y=11 8x-11=9y 8/9x-11/9=y

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome